Instructions Willow Creek Company purchased and installed carpet in its new general offices on April 30 for a total cost of $Chart of Accounts Willow Creek Company General Ledger ASSETS REVENUE 110 Cash 410 Sales 111 Petty Cash 610 Interest Revenue 1EXPENSES 510 Cost of Merchandise Sold 520 Salaries Expense 521 Advertising Expense 522 Depreciation Expense-Carpet 523 DeliveLIABILITIES 210 Accounts Payable 211 Salaries Payable 213 Sales Tax Payable 214 Interest Payable 215 Notes Payable EQUITY 310Journal PAGE 1 JOURNAL ACCOUNTING EQUATION DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY B. Record the D
Answer #1
JOURNAL ACCOUNTING EQUATION
DATE DESCRIPTION POST REF DEBIT CREDIT ASSETS LIABILITIES EQUITY
1 April 30 Carpet 123 $       35,136 $   35,136
2 April 30 Cash 110 $       35,136 $ -35,136
Dec 31 Depreciation Expense - Carpet 522 $         1,464 $ -1,464
Dec 31 Accumulated Depreciation - Carpet 124 $         1,464 $   -1,464

Depreciation = (Total Cost - Residual value) / Useful Life = (35136 - 0) / 16 = $ 2,196

Partial depreciation = 2,196 X 8 / 12 = 1,464